# DG Singular equivalence and singular locus

### v5

@@ Abstract.
For a commutative Gorenstein Noetherian ring $R$, we construct an affine scheme $X$ solely from DG singularity category $S_{dg}(R)$ of $R$ such that there is a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$. As an application, for two such rings with equivalent DG singularity categories, we prove that the singular loci in their affine schemes have the same dimension.
@@ 1. Introduction
1.1. Background
Let $A$ be an associative algebra over a base field $k$ of characteristic zero. Its singularity category $D_{sg}(A)$ is the Verdier quotient $D^{b}(A)/\mathrm{Perf}(A)$, where $\mathrm{Perf}(A)$ is the full subcategory consisting of perfect complexes over $A$. $D_{sg}(A)$ measures the smoothness of $A$ in the sense that $A$ is homologically smooth if and only if $D_{sg}(A)$ is trivial. It was first introduced by Buchweitz in the study of algebraic representations of Gorenstein rings in [5]. Moreover, he showed that $D_{sg}(A)$ is equivalent to the stable category $\underline{\mathrm{CM}}(A)$ of Cohen-Macaulay $A$-modules as triangulated categories when $A$ is Gorenstein. Later, in [21], Orlov rediscovered this notion from the perspective of algebraic geometry and mathematical physics, which has a deep relationship with Homological Mirror Symmetry.
In recent years, many mathematicians have studied singular categories from various perspective and made significant progress in this field, such as in tilting theory ([13, 23, 6] etc.), homological algebra ([5, 8, 17, 26, 27]), algebraic geometry ([28, 15, 3] etc.) and even in knot theory ([14]).
On the other hand, invariants under triangulated equivalence have been extensively studied and played an important role in the research of triangulated categories. It is widely known that the Hochschild cohomology and Hochschild homology are both invariants under the derived Morita equivalence. In [1], Armenta and Keller showed that the differential calculus of on the Hochschild cohomology and cohomology of an associative algebra is invariant under the derived Morita equivalence. Regarding the triangulated equivalence of singularity categories, Wang showed that the Gerstenhaber bracket structure on the Tate-Hochschild cohomology is invariant under the singular equivalence of Morita type with level (see [27]).
In noncommutative algebraic geometry, people focus more on invariants that arise from geometry. For example, Bondal and Orlov showed that for a projective variety $X$ with ample or anti-ample canonical bundle, its bounded derived category of coherent sheaves $D^{b}(X)$ recovers $X$ ([4]). Recently, in [11], Hua and Keller showed that the singularity category of a hypersurface algebra with isolated singularity recovers the algebra itself via the isomorphism between the zeroth Tate-Hochschild cohomology and the Tyurina algebra of this hypersurface.
@@ 1.2. Main result
In [22], Orlov proved that the completion of a variety along its singular locus determines its singularity category, up to the idempotent completion of a triangulated category. We want to answer the inverse question of Orlov’s result up to some extent.
Unfortunately, for general commutative Gorenstein Noetherian ring $R$, the information of the triangulated category $D_{sg}(R)$ is not enough to detect the geometry that we would expect. We have to consider some DG enhancement of $D_{sg}(R)$. The singularity category $D_{sg}(R)$ admits a canonical DG enhancement given by the DG quotient $S_{dg}(R):=\mathcal{D}^{b}(R)/\mathcal{P}\mathrm{erf}(R)$ ([10, §3]), where $\mathcal{D}^{b}(R)$ is the canonical DG enhancement of $D^{b}(R)$, which then induces a DG enhancement $\mathcal{P}\mathrm{erf}(R)$ of $\mathrm{Perf}(R)$. Our main result is the following.
@@
where $X_{\mathbf{p}}$ is the localization of $X$ at point $\mathbf{p}$.
@@ 1.3. Idea of the proof
The main idea of the proof of the above result is as follows.
@@
    Second, we continue to show that the zeroth Tate-Hochschild cohomology is a finitely generated module over the coordinate ring of its singular locus.
@@
    At last, we construct scheme $X$ and morphism $X\rightarrow\mathrm{Spec}(R/I)$, and then prove Theorem 1.1.
@@ 1.4. Notation and conventions
Throughout the paper, $k$ is a perfect field with characteristic zero. Unless otherwise specified, an algebra is unital over $k$, and all modules are right modules. We also assume that all commutative Noetherian rings are finitely generated over $k$ and have finite Krull dimension. For an algebra $A$, we denote by $\mathrm{mod}(A)$ the category of finitely generated $A$-modules, and by $\mathrm{proj}(A)$ its full subcategory of finitely generated projective $A$-modules. Recall that $D^{b}(A)$ is the bounded derived category of $\mathrm{mod}(A)$, and $\mathrm{Perf}(A)$ is the bounded homotopy category of complexes of $\mathrm{proj}(A)$.
@@ 2.1. Singularity category and singular equivalence
We recall the definitions of singularity category and singular equivalence for algebras.
@@
    for any $M^{\ast},N^{\ast}\in\mathrm{Ob}\big{(}D_{sg}(A)\big{)}$ and $f\in\mathrm{Hom}_{D^{b}(A)}(M^{\ast},N^{\ast})$ such that $\mathrm{Cone}(f)\in\mathrm{Ob}\big{(}\mathrm{Perf}(A)\big{)}$, let $f$ be invertible in $D_{sg}(A)$. Denote by $\mathcal{S}$ the set of morphisms satisfying the above condition in $D^{b}(A)$. Then the homomorphism space $\mathrm{Hom}_{D_{sg}(A)}(M^{*},N^{*})$ of $D_{sg}(A)$ is the localization of $\mathrm{Hom}_{D^{b}(A)}(M^{*},N^{*})$ over $\mathcal{S}$.
@@
Let $A$ and $B$ be two associative algebras. $A$ and $B$ are called singular equivalent if there is a triangle equivalence functor
@@
There are many examples of singular equivalence. For example, Chen and Sun introduced the singular equivalence of Morita type in [8]. More generally, Wang introduced singular equivalence of Morita type with level in [25].
Example 2.3.
Let $S:=k[x_{1},\cdots,x_{m}]$ for some integer $m$ and $0\neq f\in S$ be nontrivial element. Consider two algebras $S[u]/(f)$ and $S[u,v]/(f+uv)$ by adding variables $u$ and $u,v$ respectively. There are two algebra morphisms
@@
When we view $S[u]/(f)$ as an $S[u,v]/(f+uv)$-module, it admits a projective resolution as follows:
@@ 2.2. DG category and DG functor
A DG category is a $k$-linear category $\mathcal{D}$ such that the Hom-set $\mathrm{Hom}_{\mathcal{D}}(X,Y)$ consists of complexes of vector spaces, and the composition
@@
is a morphism of complexes.
A DG algebra is a DG category with one object. A homomorphism between DG algebras can be viewed as a DG functor between the corresponding DG categories.
For any DG category $\mathcal{D}$, we can construct a $k$-linear graded category $\mathrm{H}(\mathcal{D})$ given as follows:
@@
    $\mathrm{Obj}\big{(}\mathrm{H}(\mathcal{D})\big{)}=\mathrm{Obj}(\mathcal{D})$;
@@
Furthermore, let category $\mathrm{H}^{0}(\mathcal{D})$ be such that
@@
    $\mathrm{Obj}\big{(}\mathrm{H}^{0}(\mathcal{D})\big{)}=\mathrm{Obj}\big{(}\mathcal{D}\big{)}$;
@@ A DG functor $F:\mathcal{D}\rightarrow\mathcal{D}^{\prime}$ is called a quasi-equivalence if
is a quasi-isomorphism, for any $X,Y\in\mathrm{Obj}(\mathcal{D})$, and $\mathrm{H}^{0}(F)$ is an equivalence.
@@ 3.1. Generalized Tate-Hochschild complex of associative algebra
From [10, §3], we know that, for any DG category $\mathcal{A}$ and its full DG subcategory $\mathcal{B}\subset\mathcal{A}$, the DG quotient of $\mathcal{A}$ by $\mathcal{B}$ exists. On the other hand, it is well-known that both $D^{b}(A)$ and $\mathrm{Perf}(A)$ admit a canonical DG category structure. Here, we denote by $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ the DG categories corresponding to $\mathrm{Perf}(A)$ and $D^{b}(A)$ respectively. Moreover, both $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ have canonical pre-triangulated structures. Hence, by taking the DG quotient, $D_{sg}(A)$ is endowed with a DG category structure which is a DG enhancement for $D_{sg}(A)$. We denote by $S_{dg}(A)$ this DG category associated to $D_{sg}(A)$.
@@
It is well-known that the DG singular equivalence between $S_{dg}(A)$ and $S_{dg}(B)$ implies that $D_{sg}(A)\simeq D_{sg}(B)$. But, in general, singular equivalence cannot be lifted to DG singular equivalence. However, we have the following proposition for some spacial case.
@@
We continue the study of Example 2.3. In the argument of Example 2.3, we know that $M$ is a projective $S[u,v]/(f+uv)$-module, where $M:=S[u]/(f)$. Moreover, it is easy to check that $M$ is a projective $S^{op}$-module. Hence, we get a triangle functor
@@
In [17, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$ in the following theorem.
Theorem 3.6.
@@
In above proposition, the morphism $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},N^{\ast})$ in above proposition is given by composition of morphisms in $\mathcal{D}^{b}(A)$. In the meantime, we have the following.
@@
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [16, Theorem 8.1], the triangle functor
@@
induced by $\Psi_{A}$, is a triangle equivalence. In the same way, we get the equivalence
@@
If we replace $N^{\ast}$ by its projective $A$-module resolution, say, $P_{N^{*}}$, then by [19, 3.5], $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$. It follows that
@@
On the other hand, we know
@@
Now, by adjunction of functors, we obtain that
@@
where $\mathbb{D}(-):=\mathbb{R}\mathrm{Hom}_{k}(-,k)$. Note that, for any object $X$ of $\mathcal{P}\mathrm{erf}(A)$, $\mathbb{D}(\tilde{h}_{M^{\ast}})(X):=\mathbb{D}\big{(}\tilde{h}_{M^{\ast}}(X)\big{)}$. Meanwhile, we have
@@
Since $\Psi^{op}_{A}$ is a triangle equivalence, we have that $\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)^{op}}\big{(}h_{N^{\ast}},\mathbb{D}(\tilde{h}_{M^{\ast}})\big{)}$ is isomorphic to $\mathbb{R}\mathrm{Hom}_{A}\big{(}N^{*},\mathbb{D}((M^{\ast})^{\vee^{\mathbb{L}}})\big{)}$ in $D(k)$. Thus, $\mathbb{R}\mathrm{Hom}_{k}(\eta_{N^{*}}^{M^{*}},k)$ is a quasi-isomorphism. It implies that
@@
in $D(k)$. We thus completed the proof. ∎
Combining Proposition 3.7 with Lemma 3.8, we obtain that
@@
in $D(k)$. Now let $R$ be a commutative Noetherian ring, and in the above isomorphism set $A=R^{e}$ and $M^{\ast}=N^{\ast}=R$ as $R\otimes R^{op}$-module. Then we get
@@
Now, we consider the Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$ over $R^{e}$. There is a double complex:
@@
given by $\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. View this double complex as a cochain complex, and denote it by $E^{\ast}$. Also, denote by $F^{\ast}$ the cochain complex obtained from the double complex $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$. We get a complex $\mathrm{Cone}(E^{\ast}\rightarrow F^{\ast})$ and obtain the following proposition for this complex.

### v6

@@ Abstract.
For a commutative Gorenstein Noetherian ring $R$, we construct an affine scheme $X$ solely from the DG singularity category $S_{dg}(R)$ of $R$ such that there is a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$. As an application, for such two rings with equivalent DG singularity categories, we prove that singular loci in their affine schemes have the same dimension.
@@ 1. Introduction
1.1. Backgrounds
In 1986, Buchweitz introduced a quotient category of an associative algebra $A$, which is called singularity category of $A$ now [5]. The singularity category $D_{sg}(A)$ is defined as the Verdier quotient $D^{b}(A)/\mathrm{Perf}(A)$, where $\mathrm{Perf}(A)$ is the full subcategory consisting of perfect complexes in the derived category $D^{b}(A)$ of $A$. Hence, $D_{sg}(A)$ measures the smoothness of $A$ in the sense that $A$ is homologically smooth if and only if $D_{sg}(A)$ is trivial. In that note, Buchweitz showed that $D_{sg}(A)$ is equivalent to the stable category $\underline{\mathrm{CM}}(A)$ of Cohen-Macaulay $A$-modules as triangulated categories when $A$ is Gorenstein. Later Orlov rediscovered the singularity category via the perspectives of algebraic geometry and mathematical physics. Also he recover its deep relationship with Homological Mirror Symmetry [21]. Therefore singularity categories attract the interest of many mathematicians and there are many significant progress about singularity categories in various fields, such as tilting theory ([13, 23, 6] etc.), homological algebra ([5, 8, 17, 26, 27]), algebraic geometry ([28, 15, 3] etc.) and even in knot theory ([14]).
Singularity categories are triangulated categories. We know invariants under triangulated equivalence play an important role in the research of triangulated categories. For example, the Hochschild (co)homology of an associative algebra are both invariant under the derived Morita equivalence. Moreover, Armenta and Keller [1] showed that the differential calculus (consisting of Hochschild homology and cohomology endowed with several actions) of an associative algebra is invariant under the derived Morita equivalence. For singularity categories, Wang showed that the Gerstenhaber bracket structure on the Tate-Hochschild cohomology is invariant under the singular equivalence of Morita type with level (see [27]).
In noncommutative algebraic geometry, invariants from geometry are also important and interesting. Bondal and Orlov [4] showed that for a projective variety $X$ with ample or anti-ample canonical bundle, its bounded derived category of coherent sheaves $D^{b}(X)$ recovers $X$. Recently, Hua and Keller [11] showed the coordinate ring of a hypersurface with issolated singularity can be constructred via it singularity category. The authors prove the result by using an isomorphism between the zeroth Tate-Hochschild cohomology and the Tyurina algebra of the hypersurface.
@@ 1.2. Main result
In [22], Orlov proved that the completion of a variety along its singular locus determines its singularity category, up to the idempotent completion of a triangulated category. Hence there is a natural question: is it true that the singularity category of a commutative ring determines its singular locus? In this paper, We give an confirm answer of this question up to some extent.
For a commutative Gorenstein Noetherian ring $R$, $D_{sg}(R)$ is not enough to detect the geometry generally. The result in [10] motivates us to consider some DG enhancement of $D_{sg}(R)$. The singularity category $D_{sg}(R)$ admits a canonical DG enhancement given by the DG quotient $S_{dg}(R):=\mathcal{D}^{b}(R)/\mathcal{P}\mathrm{erf}(R)$ ([10, §3]), where $\mathcal{D}^{b}(R)$ is the canonical DG enhancement of $D^{b}(R)$, which then induces a DG enhancement $\mathcal{P}\mathrm{erf}(R)$ of $\mathrm{Perf}(R)$. Our main result is the following.
@@
where $X_{\mathbf{p}}$ is the localization of $X$ at point $\mathbf{p}$. Note that the scheme structure of $\mathrm{Sing}(X)$ is induced from $X$. When $X$ is an affine scheme, i.e. $X\cong\mathrm{Spec}(R)$ for some commutative Noetherian ring $R$, then
$\mathrm{Sing}(X)\cong\mathrm{Spec}(R/I)$
as scheme, where $I=\bigcap\limits_{\begin{subarray}{c}\mathbf{p}\subseteq R\\ R_{\mathbf{p}}\,is\,not\,regular\end{subarray}}\mathbf{p}$. Note that $I\cong\sqrt{I}$ as ideal of $R$ and then $\mathrm{Sing}(X)$ is reduced in $\mathrm{Spec}(R)$.
@@ 1.3. Idea of the proof
To prove the theorem, there are $3$ steps as follows
@@
    Second, we continue to show that the reduced zeroth Tate-Hochschild cohomology is a finitely generated module over the coordinate ring of its singular locus.
@@
    At last, we construct a scheme $X$ with a morphism $X\rightarrow\mathrm{Spec}(R/I)$, and prove the Theorem 1.1.
@@ 1.4. Notation and conventions
Throughout this paper, $k$ is a field with characteristic zero. In addition, commutative Noetherian rings are assumed to be finitely generated over $k$ with finite Krull dimensions. Unless specified, an associative algebra $A$ is a unital over $k$. We denote by $\mathrm{mod}(A)$ the category of finitely generated right $A$-modules and by $\mathrm{proj}(A)$ the full subcategory of $\mathrm{mod}(A)$ consisting of finitely generated projective $A$-modules. Recall that $D^{b}(A)$ is the bounded derived category of $\mathrm{mod}(A)$, and $\mathrm{Perf}(A)$ is the bounded homotopy category of complexes of $\mathrm{proj}(A)$.
@@ 2.1. Singularity category and singular equivalence
In this section, we recall the definitions of singularity category and singular equivalence for algebras.
@@
    for $M^{\ast},N^{\ast}\in\mathrm{Ob}\big{(}D_{sg}(A)\big{)}$, the set $\mathrm{Hom}_{D_{sg}(A)}(M^{*},N^{*})$ is the localization of $\mathrm{Hom}_{D^{b}(A)}(M^{*},N^{*})$ over $\mathcal{S}$, where $\mathcal{S}$ consists of $f\in\mathrm{Hom}_{D^{b}(A)}(M^{\ast},N^{\ast})$ satisfying $\mathrm{Cone}(f)\in\mathrm{Ob}\big{(}\mathrm{Perf}(A)\big{)}$.
Note that $D_{sg}(A)$ is a triangulated category [5].
@@
Let $A$ and $B$ be associative algebras, then $A$ is said to be singular equivalent to $B$ if there exists a triangulated equivalence
@@
For example, Chen and Sun introduced the singular equivalence of Morita type in [8]. More generally, Wang introduced singular equivalence of Morita type with level in [25].
Example 2.3 (Knorrer’s periodicity theorem).
Let $S:=k[x_{1},\cdots,x_{m}]$ for some integer $m$ and nontrivial element $f\in(x_{1},\cdots,x_{m})^{2}$. Considering these two algebras $S[u]/(f)$ and $S[u,v]/(f+uv)$, there exist an injective morphism
@@
$S[u]/(f)$, viewed as an $S[u,v]/(f+uv)$-module, admits a projective resolution
@@ 2.2. DG category and DG functor
A DG category is a $k$-linear category $\mathcal{D}$ such that the Hom-set $\mathrm{Hom}_{\mathcal{D}}(X,Y)$ consists of complexes of vector spaces with the associative composition
@@
is a complex morphism. definition
For example, a DG algebra is a DG category with one object. A homomorphism between DG algebras can be viewed as a DG functor between the corresponding DG categories.
For any DG category $\mathcal{D}$, there is a $k$-linear graded category $\mathrm{H}(\mathcal{D})$ given as follows:
@@
    $\mathrm{Ob}\big{(}\mathrm{H}(\mathcal{D})\big{)}=\mathrm{Ob}(\mathcal{D})$;
@@
Furthermore, denote by $\mathrm{H}^{0}(\mathcal{D})$ the category:
@@
    $\mathrm{Ob}\big{(}\mathrm{H}^{0}(\mathcal{D})\big{)}=\mathrm{Ob}\big{(}\mathcal{D}\big{)}$;
@@ A DG functor $F:\mathcal{D}\rightarrow\mathcal{D}^{\prime}$ is called a quasi-equivalence if
is a quasi-isomorphism, for any $X,Y\in\mathrm{Ob}(\mathcal{D})$, and $\mathrm{H}^{0}(F)$ is an equivalence.
@@ 3.1. Generalized Tate-Hochschild complex of associative algebra
From [10, §3], we know that, for any DG category $\mathcal{A}$ and its full DG subcategory $\mathcal{B}\subset\mathcal{A}$, the DG quotient of $\mathcal{A}$ by $\mathcal{B}$ exists. On the other hand, it is well-known that both $D^{b}(A)$ and $\mathrm{Perf}(A)$ are DG categories. Here, we denote by $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ the DG categories corresponding to $\mathrm{Perf}(A)$ and $D^{b}(A)$ respectively. Moreover, both $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ have canonical pre-triangulated structures. Hence, by taking the DG quotient, $D_{sg}(A)$ is endowed with a DG category structure which is a DG enhancement for $D_{sg}(A)$. We denote by $S_{dg}(A)$ this DG category associated to $D_{sg}(A)$.
@@
It is well-known that the DG singular equivalence between $S_{dg}(A)$ and $S_{dg}(B)$ implies that $D_{sg}(A)\simeq D_{sg}(B)$. But, in general, singular equivalence cannot be lifted to DG singular equivalence. However, we have the following proposition for some special cases.
@@
Let us back to Example 2.3. In the argument of Example 2.3, we know that $M$ is a projective $S[u,v]/(f+uv)$-module, where $M:=S[u]/(f)$. Moreover, it is easy to check that $M$ is a projective $S^{op}$-module. Hence, we get a triangle functor
@@
In [17, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$.
Theorem 3.6 ([17],Theorem 1.1).
@@
In the above proposition, the morphism $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}\rightarrow\mathrm{Hom}_{\mathcal{D}^{b}(A)}(M^{\ast},N^{\ast})$ is given by the composition of morphisms in $\mathcal{D}^{b}(A)$. In the meantime, we have the following.
@@
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [16, Theorem 8.1], the induced functor
@@
gives a triangulated equivalence. In the same way, we get the equivalence
@@
If we replace $N^{\ast}$ by its projective $A$-module resolution $P_{N^{*}}$, $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$ by the result of [19, 3.5]. It follows that
@@
On the other hand, we have
@@
Now, by the property of adjunction functors, we obtain
@@
where $\mathbb{D}(-):=\mathbb{R}\mathrm{Hom}_{k}(-,k)$. Note that for any object $X$ in $\mathcal{P}\mathrm{erf}(A)$, $\mathbb{D}(\tilde{h}_{M^{\ast}})(X):=\mathbb{D}\big{(}\tilde{h}_{M^{\ast}}(X)\big{)}$. Meanwhile, we have
@@
Since $\Psi^{op}_{A}$ gives a triangulated equivalence, we obtain that $\mathrm{Hom}_{\mathcal{P}\mathrm{erf}(A)^{op}}\big{(}h_{N^{\ast}},\mathbb{D}(\tilde{h}_{M^{\ast}})\big{)}$ is isomorphic to $\mathbb{R}\mathrm{Hom}_{A}\big{(}N^{*},\mathbb{D}((M^{\ast})^{\vee^{\mathbb{L}}})\big{)}$ in $D(k)$. Then $\mathbb{R}\mathrm{Hom}_{k}(\eta_{N^{*}}^{M^{*}},k)$ is a quasi-isomorphism, which implies
@@
in $D(k)$. Thus, we complete the proof. ∎
Combining Proposition 3.7 with Lemma 3.8, we obtain
@@
in $D(k)$. Now let $R$ be a commutative Noetherian ring, and set $A=R^{e}$ and $M^{\ast}=N^{\ast}=R$ as $R\otimes R^{op}$-module in the above isomorphism, we get
@@
Now let us consider the Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$ over $R^{e}$. There is a double complex:
@@
given by $\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. View this double complex as a cochain complex, and denote it by $E^{\ast}$. Also, denote by $F^{\ast}$ the cochain complex obtained from the double complex $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$. We get a complex $\mathrm{Cone}(E^{\ast}\rightarrow F^{\ast})$ and we obtain the following proposition.